### v2

# Recursive Inference Scaling: A Winning Path to Scalable Inference in Language and Multimodal Systems

Abstract
Recent research in language modeling reveals two scaling effects: the well-known improvement from increased training compute, and a lesser-known boost from applying more sophisticated or computationally intensive inference methods. Inspired by recent findings on the fractal geometry of language, we introduce Recursive INference Scaling (RINS) as a complementary, plug-in recipe for scaling inference time. For a given fixed model architecture and training compute budget, RINS substantially improves language modeling performance. It also generalizes beyond pure language tasks, delivering gains in multimodal systems, including a +2% improvement in 0-shot ImageNet accuracy for SigLIP-B/16. Additionally, by deriving data scaling laws, we show that RINS improves both the asymptotic performance limits and the scaling exponents. These advantages are maintained even when compared to state-of-the-art recursive techniques like the “repeat-all-over” (RAO) strategy in Mobile LLM. Finally, stochastic RINS not only can enhance performance further but also provides the flexibility to optionally forgo increased inference computation at test time with minimal performance degradation.
Machine Learning, ICML
1 Introduction
There has been a proliferation of research in recent years pointing to the pivotal role of scale, and how its benefits could be predicted empirically (Mukherjee et al., 2003; Hestness et al., 2017; Johnson et al., 2018; Rosenfeld et al., 2019; Kaplan et al., 2020; Ghorbani et al., 2021; Alabdulmohsin et al., 2022; Bansal et al., 2022; Zhai et al., 2022). Generally, the performance of deep neural networks $f(x)$ (such as its error rate or log-perplexity) often follows a power law $f(x)\sim\beta x^{-c}+\varepsilon$ as one varies a dimension of interest $x$, such as the size of the training data or model parameters. These “scaling laws,” as they are known in the literature, have been used, among others, to determine the training data size needed to achieve a specified level of accuracy (Cho et al., 2015; Beleites et al., 2013; Figueroa et al., 2012) and to optimize the model architecture (Kaplan et al., 2020; Hoffmann et al., 2022; Alabdulmohsin et al., 2024b). They have also been justified theoretically (Bahri et al., 2021; Hutter, 2021; Sharma & Kaplan, 2022).
Figure 1: left: Illustration of Recursive INference Scaling (RINS) compared to a standard forward pass. In RINS, the model $f:\mathcal{X}\to\mathcal{Y}$ is split into two parts: the first block $f_{A}:\mathcal{X}\to\mathcal{X}$ is applied iteratively to its own output for $r$ iterations before passing the output to the second block. right: Illustrative examples of models with different signatures and degrees. From top to bottom: (1) Baseline (Signature: AB, Degree: 1), a standard feedforward architecture with no recursion. (2) repeat-all-over (RA) (Liu et al., 2024), where the entire model is recursively applied on its output. When recursion is done twice, it has a signature of ABAB. (3) RINS with signature A3B. (4) (A3B)2 whose degree is 2, in which the same parameter sharing signature is applied on each of the two blocks A and B.
Besides this conventional approach of scaling the training compute, the impact of increased inference compute on model performance has emerged as another key scaling dimension. Whereas OpenAI’s GPT-o3 has prominently demonstrated this recently (OpenAI, 2024), earlier work had already explored this direction. For example, chain-of-thought (CoT) prompting has shown that eliciting longer inference paths through additional token generation could improve the reasoning capabilities of language models (Wei et al., 2024). This is similar to the demonstrated success of critiquing before evaluating (Ankner et al., 2024). Also, AlphaCode (Li et al., 2022) and Codex (Chen et al., 2021) generate multiple samples during inference in order to enhance code generation performance. Remarkably,  Brown et al. (2024) shows that the benefit of sampling multiple solutions in tasks such as mathematics and coding—when measured by coverage—holds for up to four orders of magnitude of inference compute. These approaches suggest that, like training compute, inference compute follows systematic scaling patterns that can be leveraged to improve models. Refer to the survey (Welleck et al., 2024) for more details.
Recently, it has also been noted that language exhibits a “self-similar” (fractal) nature, meaning that similar patterns repeat across different scales of its representation, from individual words to entire sentences and paragraphs (Alabdulmohsin et al., 2024a). Inspired by this finding, we examine if model recursion, which can be interpreted as a particular form of scale-invariant decoding, offers a complementary approach for scaling inference time in language models. To this end, we examine an extensive set of parameter-sharing strategies and, indeed, identify the best to be a special form of recursion, which we term Recursive INference Scaling (RINS). We show that RINS yields significant performance gains compared to other parameter-sharing methods, even when controlling for model size and training compute.
RINS builds upon the concept of model recursion but recasts it as a powerful inference-time scaling strategy. It leverages a simple yet profound idea: use your existing architecture and training compute budget as is, but exploit the self-similar structure of language by recursively applying an early portion of your network to refine its output. In turn, this simple strategy improves performance significantly, with the primary cost being a scaled inference time.
Recursion has shown promise in language modeling, with recent work by Liu et al. (2024) demonstrating that recursive architectures outperform similarly sized vanilla models trained on the same number of tokens. Notably, Liu et al. (2024) finds that a simple “repeat-all-over” (RAO) approach, where the entire model is recursively applied, surpasses more complex parameter-sharing strategies in language.
However, while Liu et al. (2024) demonstrate the sample efficiency of recursive architectures, in which models are compared when trained on the same number of tokens, their analysis does not explicitly account for the increased computational cost of recursive operations during training. Hence, it remains unclear whether the performance gains observed in prior work come from the inherent advantages of model recursion or simply from having increased the training compute. Indeed, our findings suggest that for moderately sized models (over 1 billion parameters), the performance gains of the “repeat-all-over” (RAO) strategy can be matched by extending the training of a standard model to consume an equivalent compute. RINS, by contrast, significantly outperforms all other baselines on a compute- and parameter-matched setup, including when scaling inference by increasing the context length (see Figure 2).
Crucially, a stochastic variant of RINS, in which the number of recursion rounds is sampled at training time from a binomial distribution, not only can enhance performance further, such as in multimodal systems, but also provides the flexibility to optionally forgo increased inference computation at test time with minimal performance degradation. See Section 4 for details.
Efficiency.
We conduct our experiments mostly on small models, which are typically intended for deployment environments with stringent memory limitations (Liu et al., 2024). Given the direct relation between a model’s memory footprint and its parameter count (e.g. a 1 billion parameter model with 16-bit floating-point precision requires 2GB of DRAM), the ability to enhance accuracy while maintaining a fixed parameter count is highly desirable. RINS achieves this by enabling significant performance gains without increasing parameter count for the same training compute.
Statement of Contribution. In summary, we introduce Recursive INference Scaling (RINS), a complementary plug-in method for scaling inference time. We:
  1. 1.
    propose a taxonomy of parameter-sharing architectures, empirically evaluating their effectiveness across multiple domains. Our comprehensive analysis identifies RINS as a powerful approach, outperforming methods like RAO used in Mobile LLM (Liu et al., 2024), and scaling inference by increasing the sequence length.
  2. 2.
    argue that RINS’ performance gain likely stems from the self-similar (fractal) nature of language, as a similar analysis in vision yields minimal improvements.
  3. 3.
    use RINS to extend inference scaling beyond language to multimodal systems that incorporate language in their processing, such as contrastive models. In particular, our SigLIP-RINS-B/16 outperforms the popular SigLIP-B/16 (Zhai et al., 2023) by a wide margin; e.g. improving 0-shot accuracy in ImageNet from 77.3% to 79.6% and CIFAR100 from 70.3% to 80.7%.
  4. 4.
    derive data scaling laws for RINS, revealing improvements in both the asymptotic performance limit and convergence speed (i.e. scaling exponent).
  5. 5.
    show that stochastic RINS can enhance performance even further, such as in multimodal systems, while offering the option to revert to non-recursive inference at test time with minimal performance degradation.
2 Recursive Inference Scaling
2.1 Overview
Before describing how RINS works, we briefly formalize definitions. Let $\mathcal{X}$ be a fixed domain, often the space of sequences of soft tokens of embedding dimension $d$. Let $\mathbb{L}=\{l_{1},l_{2},...,l_{n}\}$ be a fixed set of $n$ unique blocks, where each block $l_{k}:\mathcal{X}\to\mathcal{X}$ is a function mapping from the input space $\mathcal{X}$ to the same output space $\mathcal{X}$. By “unique” here we simply imply that such blocks (which typically comprise of multiple layers each) are not constrained to share the same parameters. Let $\mathcal{G}(\mathbb{L})$ be the space of all possible computation graphs representing neural network architectures that can be constructed by composing blocks from the set $\mathbb{L}$, while $f\in\mathcal{G}(\mathbb{L})$ be one specific architecture.
Figure 1 (right) illustrates some examples for the case when $|\mathbb{L}|=2$. For instance, one can repeatedly apply the entire model, as in the “repeat-all-over” (RAO) approach in Mobile LLM (Liu et al., 2024), or recursively apply a strict subset of the architecture, such as a single a block within the model. The choice of arrangement of blocks can significantly impact the model’s performance and efficiency.
Formally, let $C(f)$ be the actual computational cost (in FLOPs) of training $f\in\mathcal{G}(\mathbb{L})$ on a dataset sampled i.i.d. from distribution $\mathcal{D}$, considering only the forward pass. Also, $\mathcal{L}(f)$ is a performance metric of interest (e.g., validation loss or error rate) for model $f$. We assume that a lower value of $\mathcal{L}$ indicates better performance.
Definition 2.1.
For a fixed set of unique blocks $\mathbb{L}$ and a training compute budget $c$, a recursive architecture $f^{\star}\in\mathcal{G}(\mathbb{L})$ is considered “better” than another architecture $f\in\mathcal{G}(\mathbb{L})$ if $C(f^{\star})\leq C(f)\leq c$ and $\mathbb{E}[\mathcal{L}(f^{\star})]\leq\mathbb{E}[\mathcal{L}(f)]$.
In other words, we search for the architecture $f^{\star}$, constructed only from the set of blocks $\mathbb{L}$, that minimizes the loss under the constraint of a bounded training compute $c$.
Model recursion offers a simple, plug-in approach for scaling inference time. By applying some layers iteratively to their own output, we effectively increase the computational path length during inference without altering the underlying model architecture. This allows us to exploit the benefits of increased inference compute. Importantly, it is complementary to other techniques like chain-of-thought (CoT) prompting (Wei et al., 2024), which achieve performance gains by eliciting longer generation sequences, and repeated sampling (Li et al., 2022; Chen et al., 2021).
Figure 2: Language models are trained on 200B tokens. The $x$-axis is the training cost in units of layer $\times$ step. Notably, the performance advantage of RINS increases with increased computational budget. The long-sequence baseline, using a context length of 1,536 tokens, exhibits lower performance due to processing fewer examples to maintain the same FLOPs count. See Figure 4 for results with extended training duration and Figure 5 for results with larger (1B parameter) models, further demonstrating the benefits of RINS.
Taxonomy.
As discussed in Section 1, language has been shown to exhibit self-similarity, meaning that similar patterns repeat across multiple levels of granularity, from the structure of individual sentences to the composition of entire texts (Alabdulmohsin et al., 2024a). This observation suggests that recursive (scale-invariant) decoding could be particularly beneficial for language processing.
To determine the extent to which this is true, we systematically explore a vast space of parameter sharing strategies. We use the following taxonomy based on two key concepts: (1) signature and (2) degree. The “signature” of an architecture describes how parameter sharing is applied across different blocks of the model. For example, an architecture with the signature AB2C indicates that the model is partitioned into three unique blocks (A, B, and C) of equal size. The processing flow in this architecture would be: apply block A on the input, apply block B to A’s output, apply block B again to its own output (hence the exponent), and finally apply C to the output of B. Note that while the signature AB2C has the same parameter count as a simpler ABC architecture, it incurs approximately 33% more compute due to the repeated application of block B. This highlights the trade-off between computational cost and performance using recursion that we will carefully study in this work.
On the other hand, “degree” specifies whether recursion is applied only at the level of entire blocks, or if it extends to individual components within blocks, as illustrated in Figure 1 (bottom). A degree of 2, for example, means that each of the blocks A, B, and C in our example have themselves the same recursive pattern used in the overall model. This adds another dimension to our taxonomy, allowing for a finer-grained classification of recursive architectures. Note that degree makes notation concise, but is not necessary; e.g. (ABB)2 is equivalent to ABBCDDCDD. By systematically varying both signature and degree, we can comprehensively explore the space of recursive models and identify optimal configurations. Figure 3 provides a detailed pseudocode.
With this framework, we now state our main investigation:
Which family of architectures (i.e. signatures and degrees) lead to better performance according to Definition 2.1 under fixed compute budget?
To reiterate, this is a non-trivial question because it is conceivable for a non-recursive model to outperform all others given that it sees more training examples, utilizing the compute budget without resorting to recursion. For instance, increasing the context length in language modeling can be inferior to training on more examples within a limited compute budget. We present examples of this in Section 3.
Nevertheless, our analysis reveals that Recursive INference Scaling (RINS) emerges as a clear winner. Its consistent superiority suggests that it captures a fundamental principle for efficiently processing language. We hypothesize that this is due to the self-similar geometry of language.
Definition 2.2.
RINS corresponds to architectures with degree 1 and signature ArB, for some integer $r>1$.
In other words, RINS partitions a given model into two equally-sized blocks A and B. For example, A would correspond to the first 6 layers while B would correspond to the second 6 layers in a 12-layer neural network (ignoring the embedding layer and classifier head). The first block A is recursively applied on its own output $r$ times before passing the output to B. See Figure 1 for illustration and Figure 3 for a detailed pseudocode. In Section 6, we show how the optimal value of $r$ depends on the allocated compute budget.
Main Claim.
Our claim can be summarized as follows. Once a language model is chosen and the training compute budget is planned, one should enable RINS, which does not change the model size, and train the new RINS-enabled model for the same amount of compute. This is achieved by partitioning the original model into two equally-sized blocks and applying recursion to the first block alone. Thus, neither model size nor training compute is affected. Our empirical results demonstrate that RINS-enabled models will consistently outperform baseline models. In addition, stochastic RINS with some positive skip probability $p_{s}>0$ (see Section 4) offers the option of forgoing scaling inference at test time with little performance degradation.
3 Experimental Results
In this section, we study the impact of various parameter-sharing strategies in language modeling, following the taxonomy introduced in Section 2. We show how RINS emerges as a clear winner. All experiments are carried out using the Big Vision codebase (Beyer et al., 2022).
Metric Architecture
BL RAO RINS
Winogrande 53.3% 53.0% 53.3%
OpenBookQA 37.7% 38.3% 38.9%
BoolQ 47.7% 41.6% 52.5%
PIQA 54.5% 68.6% 69.5%
SIQA 39.4% 40.2% 41.3%
Table 1: Zero-shot evaluation in downstream common sense reasoning tasks. All models are 600M parameters, pretrained on the compute-equivalent of 500B tokens in the baseline (BL). The best signatures for RAO and RINS in Figure 4 are used in this evaluation. See Appendix C for details on the evaluation setup.
Setup.
First, we train a decoder-only transformer language model (Vaswani et al., 2017) with relative position embeddings and sequence packing. We use C4/English tokenizer with a vocabulary size of 32K. The model is trained on a mixture of C4 (Raffel et al., 2020) and SlimPajama (Soboleva et al., 2023) with equal weight using a batch size 1,024 and context length 1,024. The non-recursive baseline is trained for 200K steps, which amounts to 200B training tokens. Other recursive models are trained on fewer tokens in order to match the same total compute FLOPs.
The optimizer is Adam (Kingma & Ba, 2014) with learning rate $5\times 10^{-4}$ and weight decay $5\times 10^{-5}$, using an inverse square root decay schedule with 5K warm-up and cool-down steps. This learning rate was chosen by sweeping across the values $\mathrm{lr}\in\{10^{-3},5\times 10^{-4},10^{-4},5\times 10^{-5}\}$ with $\mathrm{wd}=\mathrm{lr}/10$. The reported value yielded the best performance for the non-recursive baseline. The full training configuration along with a list of all the architectures we sweep across (i.e. signature and degree) can be found in Appendix A. Each model is trained on $16\times 16$ TPUv5 chips for approximately 2.2K core-hours. Models that failed OOM were excluded. Overall, we train 59 models of two sizes: 300M and 600M parameters. In all models, the embedding dimension is 1,536 and the MLP dimension is 6,144. Later, we present results using models with 1 billion parameters.
Results.
Figure 2 shows that RINS outperforms all other approaches. Importantly, the optimal number of recursion rounds in RINS (e.g. whether to use A2B or A3B or more recursions) depends on the allocated training compute budget. This is seen in Figure 2 in the fact that the non-recursive baseline AB (a special case of RINS with $r=1$) initially outperforms all other models, before its performance saturates and recursive models begin to outperform it for the same training compute FLOPs and parameter count. We study the relation between the optimal number of recursion rounds and compute later in Section 6. In addition, RINS outperforms scaling inference time by only increasing the sequence (context) length (green vs. red curves in Figure 2).
Longer Training Duration.
One observation in Figure 2 is that the performance gap seems to widen in favor of RINS as more training compute is allocated. To verify this, we resume training the 600M-parameter models on 500B training tokens. Here, we only compare RINS with the non-recursive baseline and RAO, since RAO was identified in prior works to outperform other parameter-sharing strategies (Liu et al., 2024). Figure 4 shows that the performance gap continues to increase in favor of RINS. This improvement also persists downstream, as shown in Table 1.
4 Stochastic Recursive Inference Scaling
class RecursiveBlock():
config: Dict # single block config
signature: str = "a"
degree: int = 1
p_skip: Tuple[float, ...]] # skip prob for blocks
def __call__(self, x):
"""Call model on input x."""
if degree == 1:
blocks = {c: SingleBlock(config=self.config
) for c in set(self.signature)}
else:
blocks = {c: RecursiveBlock(
config=self.config, signature=self.signature,
degree=self.degree - 1, p_skip=self.p_skip
) for c in set(self.signature)}
inputs = x
# forward pass
for i in range(len(self.signature)):
c = self.signature[i]
choice = random.uniform()
if self.degree == 1: # stochastic RINS
if choice > self.p_skip[i]:
x = blocks[c](x) # else skip
else:
x = blocks[c](x)
return x
Figure 3: Numpy-like syntax for models with a fixed signature and degree. When no stochastic depth is applied, we have $p_{s}=0$. In RINS, we expand $p_{s}$ into a tuple of the form $(0,p_{s},p_{s},\ldots,p_{s},0)$, where the first and last entries are zero to guarantee they are executed, which is equivalent to sampling the number of recursion rounds from a binomial distribution as described in Section 4.
Figure 4: Performance of RINS (A2B, A3B, A4B) vs. RAO (A2, A3, A4) with increasing compute on 600M-parameter models. All models have the same size. The performance advantage of RINS grows with the computational budget. The long-sequence baseline is not shown since it underperforms other methods in Figure 2.
Next, we investigate the effect of stochastically dropping blocks during training, inspired by the regularization technique of stochastic depth (Huang et al., 2016). Our primary goal is to determine whether this approach can further enhance the performance of RINS while simultaneously offering the flexibility of reverting to non-recursive inference without significant degradation in model quality.
To recall, RINS has the signature ArB for some $r>1$. To implement stochastic RINS, we introduce a skip probability $p_{s}\in[0,1)$ and sample during training the number of recursion rounds in each step to be $1+\eta$, where $\eta$ is a binomial random variable with probability of success $1-p_{s}$ and number of trials $r-1$. Thus, block A is always executed at least once. During inference, we are free to choose how to scale inference compute by setting the value of $r\geq 1$. See the detailed pseudocode in Figure 3.
Here, we train bigger models with 1 billion parameters. We use an embedding dimension 2,048 and MLP dimension 8,192. All models have 18 decoder blocks. We train for 500K steps and compare RINS with signature A3B against the non-recursive baseline.
Figure 5 summarizes the advantages of stochastic RINS. Notably, we observe that as $p_{s}>0$ increases, stochastic RINS mitigates the performance degradation incurred when scaling is not applied at inference time, while still offering big gains when inference time is scaled. Not surprisingly, though, scaling inference time is less effective when $p_{s}$ increases, suggesting a tradeoff between flexibility at inference time and the magnitude of potential gains from scaling. As shown in Figure 6, similar conclusions hold even in the asymptotic (infinite-compute) regime, under the assumption that loss follows a power law relation (Kaplan et al., 2020).
Figure 5: Performance of stochastic RINS (A3B) with varying inference costs for 1B parameter LMs. The $x$-axis represents the training compute cost. The legend indicates the inference cost of each stochastic RINS configuration relative to the baseline; e.g. $1.5x$ denotes 50% increase in inference cost. For $p_{s}=0$, RINS@1x is significantly worse, with perplexity scores $>3$. As expected, RINS converges in performance to the baseline as $p_{s}\to 1$.
Figure 6: Asymptotic performance of 1B-parameter LMs, evaluated in C4 (left) and SlimPajama (right). Dotted line is for baseline. We observe a tradeoff between inference flexibility (gap between 1x & 2x) and magnitude of gain from inference scaling (2x result).
5 Other Modalities
5.1 Vision
As previously discussed, the performance gains in RINS are consistent with the self-similar nature of language. By performing a recursive, scale-invariant decoding, RINS introduces an inductive bias that encourages the model to recognize and exploit recurring patterns at different scales (see Appendix D for further discussion). To test if this is likely the source of its advantage, we conduct a similar empirical evaluation in vision, a domain lacking self-similarity.
Setup. We train an encoder-only vision transformer ViT-B/16 (Dosovitskiy et al., 2021) on ImageNet-ILSRCV2012 (Deng et al., 2009). The non-recursive baseline model is trained for either 300 or 1,000 epochs using a batch size 1,024, while recursive models are trained on fewer epochs to match the same total training compute FLOPs. We apply MixUp (probability 0.2) during training (Zhang et al., 2018) and use learned position embedding. The optimizer is Adam where we tune the learning rate for each architecture in the set $\mathrm{lr}\in\{10^{-3},7\times 10^{-4},3\times 10^{-4},10^{-4}\}$ with weight decay $\mathrm{wd}=\mathrm{lr}/10$, on a small validation split. We use a cosine learning rate schedule with 10K warm-up steps. Images are resized to $224\times 224$ and $16\times 16$ patch sizes are used. The full training configuration is in Appendix A.
Results. As presented in Table 2, parameter-sharing techniques, including RINS, do not confer any advantage in supervised image classification. The non-recursive baseline, when trained on longer sequence lengths (i.e., higher image resolution) to match the inference cost of recursive architectures, surpasses all other methods. This starkly differs from the results observed in language modeling, where RINS provides significant gains even when compared against models that scale inference by increasing the sequence length.
Architecture Val ReaL v2 Avg
300 epochs
(ABB)2 75.2 81.4 62.7 73.1
A@336 75.7 81.0 62.3 73.0
AAB 75.2 80.5 61.4 72.4
ABBC 75.1 80.2 61.3 72.2
A@224 74.9 80.1 61.3 72.1
1,000 epochs
A@336 77.6 82.7 64.6 75.0
ABBC 76.3 81.6 63.2 73.7
AABC 76.0 81.4 62.4 73.2
AA 75.8 81.4 62.4 73.3
AAA 75.7 80.8 62.0 72.8
Table 2: Performance of the top 5 architectures on ILSRCV2012 classification. The table compares the performance of recursive architectures with a baseline trained 224- and 336-resolution images. The baseline A@336, trained on higher-resolution to match the inference cost of the recursive models, outperforms all parameter-sharing architectures. We use the notation $(\mathrm{signature})_{\mathrm{degree}}$.
5.2 Contrastive Mutlimodal Systems
Setup.
Finally, we also study the impact of RINS in vision-language pretraining, motivated by the fact that such models also process natural language. We pretrain SigLIP-B/16 models (Zhai et al., 2023), which are contrastive models trained using the sigmoid loss on English-language image–text pairs. We follow (Zhai et al., 2023) in most aspects. Images are resized to $256\times 256$ and we use $16\times 16$ patch sizes. Texts, on the other hand, are tokenized using C4 tokenizer (Raffel et al., 2020) with a vocabulary size of 32K, and we keep a maximum of 64 tokens. The optimizer is Adam with learning rate $10^{-3}$ and weight decay $10^{-4}$, using an inverse square root decay schedule with 5K warm-up and cool-down steps. For the baseline, we use SigLIP-B/16 pretrained on 10B training examples. Again, recursive models are trained on fewer examples to match the total training compute cost. Due to the amount of compute involved in these experiments, we only compare the non-recursive baseline (with signature AB) against RAO (with signature ABAB) and RINS (signature A2B) with degree 1.
Metric Architecture
AB Long-Seq ABAB AAB
Zero-shot classification
ImageNet 73.4 73.0 72.7 74.1
CIAFR100 68.9 63.6 65.3 72.2
Pet 90.4 90.0 90.7 90.0
Retrieval
COCO img2txt@1 62.7 62.0 61.1 62.3
COCO img2txt@5 84.8 84.1 82.9 84.1
COCO img2txt@10 90.7 90.4 89.5 90.2
COCO txt2img@1 44.6 44.2 43.2 45.1
COCO txt2img@5 69.6 69.4 68.4 70.0
COCO txt2img@10 78.8 78.7 77.5 79.0
Flickr img2txt@1 89.6 87.2 87.9 88.9
Flickr img2txt@5 98.0 97.8 97.5 98.5
Flickr img2txt@10 99.1 98.9 98.9 99.3
Flickr txt2img@1 75.1 73.7 73.2 74.3
Flickr txt2img@5 92.3 92.1 91.5 92.4
Flickr txt2img@10 95.6 95.6 95.6 95.8
Table 3: Performance of SigLIP-B/16 on various datasets. Results are shown for ImageNet (Deng et al., 2009), CIFAR100 (Krizhevsky, 2009), Pet (Parkhi et al., 2012), COCO (Chen et al., 2015), and Flickr (Young et al., 2014). All models are identical in size to SigLIP-B/16 and have the same training compute FLOPs. Long-Seq is SigLIP-B/16 trained on higher resolution of 280 and text length 80 (25% increase in sequence length $\rightarrow$ 50% increase in inference cost compared to baseline, similar to AAB). Using Wilcoxon signed rank test (Wilcoxon, 1992), we obtain $p=0.003$ so the evidence in favor of RINS is statistically significant at the 99% confidence level.
Results. Table 3 shows that RINS (with signature A2B) outperforms the non-recursive baseline, including with long sequence length, and RAO in zero-shot and retrieval evaluations. Of importance is the impact in ImageNet 0-shot classification, where we see an improvement of about $+0.7\%$.
Overtraining. Next, we demonstrate that RINS can substantially advance state-of-the-art results in multimodal systems for a given model size. We train a recursive variant of SigLIP-B/16, denoted SigLIP-RINS-B/16, using the signature A3B. In light of the findings in Section 6, we increase the number of recursions here given the increase in compute.
We adhere to the training protocol outlined above, with the exception that SigLIP-RINS-B/16 is now trained on 40B examples, matching the training data scale of the widely-used, publicly available SigLIP-B/16 checkpoint. Note that both models have an identical size. Moreover, following Pouget et al. (2024), we utilize a training mixture comprising both English and multilingual data to enhance cultural diversity, and report the cultural metrics recommended in Pouget et al. (2024) as well as multilinguality evaluations using XM3600 dataset (Thapliyal et al., 2022). So, to ensure an appropriate comparison of results, we re-train SigLIP-B/16 on 40B examples from the same data mixture. The primary datasets for cultural diversity evaluation are Dollar Street (Rojas et al., 2022), GeoDE (Ramaswamy et al., 2024), and Google Landmarks Dataset v2 (GLDv2) (Weyand et al., 2020). We use the Gemma tokenizer (Mesnard et al., 2024).
As shown in Table 4, SigLIP-RINS-B/16 significantly outperforms the standard SigLIP-B/16 across all benchmarks. In fact, stochastic RINS with skip probability $p_{s}=\frac{1}{4}$ improves results even further. Crucially, these results demonstrate that RINS offers a fundamental advantage in multimodal learning that are not replicated by simply overtraining a non-recursive counterpart.
Metric SigLIP-B/16 SigLIP-RINS-B/16
(370M params) (370M params)
$p_{s}=0$ $\frac{1}{4}$ $\frac{1}{2}$
Zero-shot classification
ImageNet 77.3 79.0 79.6 79.2
CIAFR100 70.3 78.5 80.7 81.7
Pet 92.6 94.8 94.4 92.7
Cultural Diversity
GeoLoc:Dollar Street 17.6 19.7 19.2 19.3
GeoLoc:GeoDE-Country 22.5 23.3 24.7 23.9
GeoLoc:GeoDE-Region 36.1 37.7 39.7 38.7
Dollar Street 51.5 52.9 53.1 53.1
GeoDE 93.1 93.7 94.3 94.2
GLDv2 51.6 53.7 52.4 52.5
Multilinguality
XM3600 img2txt@1 48.4 53.5 52.6 51.8
XM3600 img2txt@5 68.4 72.6 72.0 70.9
XM3600 img2txt@10 74.4 78.1 77.5 76.5
XM3600 txt2img@1 39.5 43.0 43.1 41.9
XM3600 txt2img@5 59.5 63.0 63.1 61.6
XM3600 txt2img@10 66.1 69.3 69.3 68.0
Retrieval
COCO img2txt@1 67.6 69.4 70.0 69.5
COCO img2txt@5 87.2 88.6 88.9 88.3
COCO img2txt@10 92.6 93.2 93.5 93.2
COCO txt2img@1 50.3 51.5 52.4 52.0
COCO txt2img@5 74.7 75.2 76.0 75.7
COCO txt2img@10 82.6 83.0 83.5 83.2
Flickr img2txt@1 91.9 92.9 93.5 92.4
Flickr img2txt@5 99.3 98.7 99.5 98.7
Flickr img2txt@10 99.6 99.5 99.7 99.3
Flickr txt2img@1 80.1 80.7 81.4 80.5
Flickr txt2img@5 94.6 94.3 95.4 95.0
Flickr txt2img@10 97.1 97.0 97.4 97.0
Table 4: Performance of multilingual SigLIP models on various datasets under an overtraining regime, where training compute is not a constraint. All models are identical in size to SigLIP-B/16. As shown in the rightmost columns, stochastic RINS ($p_{s}>0$) outperforms the other architectures. Detailed results on multilinguality are provided in Appendix B.
6 Further Analysis
6.1 Data scaling laws
Next, we investigate the influence of Recursive INference Scaling (RINS) on the data scaling behavior of language models. Specifically, we fit a power law of the form $\varepsilon(x)=\beta x^{-c}+\varepsilon_{\infty}$, to the log-perplexity loss $\varepsilon(x)$ as a function of the training FLOPs $x$. This allows us to analyze the impact of RINS on both the scaling exponent $c$ and the asymptotic performance limit $\varepsilon_{\infty}$, revealing whether the performance gains in RINS stem from an improved scaling exponent, a lower asymptotic limit, or a combination of both. We use the 600M-parameter language models in Figure 4 whose signature is ArB, for $r\in\{1,2,3,4\}$, where $r=1$ corresponds to the non-recursive baseline. As shown in Figure 7, RINS improves both the scaling exponent and the asymptotic limit. The improvement in the asymptotic limit provides further evidence that the performance gap in favor of RINS is not closed by overtraining non-recursive models.
Figure 7: top: The scaling exponent $c$ in models with signature ArB, where $r=1$ corresponds to the non-recursive baseline. RINS ($r>1$) improves scaling exponents. bottom: Asymptotic log-perplexity $\varepsilon_{\infty}$ is plotted against $r$. RINS improves $\varepsilon_{\infty}$, indicating that overtraining the baseline cannot match its performance gains.
6.2 Optimal Number of Recursion Rounds
We observe from the scaling parameters in the previous section that if the performance of RINS is modeled as a power law $f_{r}(x)=\beta_{r}x^{-c_{r}}+\varepsilon_{r}$, then $c_{r}$ increases with $r$ while $\varepsilon_{r}$ decreases. Furthermore, the coefficient $\beta_{r}$ also increases with $r$. This implies that while scaling inference by increasing $r$ might initially exhibit a higher loss due to the larger $\beta_{r}$, its faster convergence ($c_{r}$) and lower asymptotic limit ($\varepsilon_{r}$) will eventually lead to superior performance. In other words, using the higher recursion level is advantageous eventually, which is consistent with the experimental results.
Figure 8: Optimal number of recursions $r$ in RINS is plotted vs. training compute for 300M-, 600M-, and 1B-parameter LMs.
To observe this more explicitly, we train language models with four signatures ArB: $r\in\{1,2,3,4\}$. Then, we plot the optimal value of $r$ against training compute. As shown from the results in Figure 8, the optimal value of $r$ monotonically increases with training compute, in agreement with earlier results. Also, smaller models benefit more from RINS.
7 Discussion and Related Works
The premise that extending inference time enhances the quality of language model outputs finds support in cognitive science. Human deliberation, often associated with System 2 thinking, is linked to improved decision-making capabilities (Lawson et al., 2020). Mirroring this, numerous computational strategies have been developed to scale inference within Large Language Models (LLMs). Prompt-based methods like Chain-of-Thought (CoT) (Wei et al., 2024) and critique-then-generate approaches (Ankner et al., 2024) are prominent examples. Other iterative refinement methods like ReAct (Yao et al., 2023), Self-Refine (Madaan et al., 2024), and Reflexion, which incorporate feedback and reflection, also improve inference quality. Even simpler strategies, capitalizing on the stochastic nature of LLM decoding that generate multiple responses and select among them, such as self-consistency (Wang et al., 2023), have shown efficacy that scales well across multiple orders of magnitude of inference compute (Brown et al., 2024).
However, the assumption of monotonic improvement with increased inference calls is not guaranteed to hold. As argued by Chen et al. (Chen et al., 2024), repeated sampling may lead to convergence towards the most probable, but not necessarily the optimal, solution. This is particularly pertinent for challenging instances where the probability of a correct answer is below chance (i.e., $<0.5$).
In this work, we introduce a complementary approach, called Recursive INference Scaling (RINS), which can be integrated with other techniques.. RINS identifies a specific form of model recursion as an effective inference scaling strategy, demonstrating significant performance gains over various other recursive architectures, including the Repeat All Over (RAO) strategy identified by Liu et al. (2024) as state-of-the-art for mobile LLMs.
Recursion is a form of parameter sharing, a technique that has been explored for a while. Illustrative examples include Universal Transformers (Dehghani et al., 2019) and ALBERT (Lan et al., 2020). Despite their innovative design, their scaling exponents were smaller than in the vanilla transformer (Tay et al., 2022a) so they failed to subsume it (Tay et al., 2022b). RINS, by contrast, improves both the scaling exponents and asymptotic performance limits.
We hypothesize that the efficacy of RINS in language tasks stems from its exploitation of language’s self-similarity (Alabdulmohsin et al., 2024a), since the recursive block in RINS is a form of scale-invariant decoding (see Appendix D for further discussion). One evidence for this is the fact that RINS does not help in the pure vision domain, but it can still improve performance in multimodal systems that also process language, such as contrastive models. We obtain significant performance gains in SigLIP, for example.
Impact Statement
This paper presents work whose goal is to advance the field of artificial intelligence, specifically in the area of language modeling and multimodal systems. There are many potential societal consequences of advancing AI, both positive, such as improved accessibility to high-quality healthcare and education, and negative, if such systems are misused. Our intention with RINS is to contribute to the ongoing effort to improve the performance and capabilities of these models. While we recognize the general ethical considerations that accompany the advancement of language models and AI, we do not feel that this particular work raises unique or unaddressed ethical concerns beyond the established considerations within the field.
Acknowledgement
We thank Michael Tschannen and Jeremiah Harmsen for their insightful reviews and suggestions, and Google DeepMind at large for providing a supportive research environment.
Appendix A Training Configurations
Appendix B Multilinguality Detailed Results
Language Image-to-Text Retrieval @ 1 Text-to-Image Retrieval @ 1
SigLIP-B/16 SigLIP-RINS-B/16 SigLIP-B/16 SigLIP-RINS-B/16
$p_{s}=0$ $p_{s}=\frac{1}{4}$ $p_{s}=\frac{1}{2}$ $p_{s}=0$ $p_{s}=\frac{1}{4}$ $p_{s}=\frac{1}{2}$
ar 54.4 58.8 59.0 58.2 42.2 45.8 46.7 45.3
bn 7.0 11.2 10.3 8.7 4.8 7.2 6.8 5.1
cs 51.1 55.8 55.8 54.5 40.5 45.0 44.6 43.0
da 61.3 68.8 68.8 65.2 45.8 51.6 52.6 49.8
de 81.5 84.1 83.3 83.3 69.3 72.2 72.4 71.5
el 38.8 45.8 44.9 42.9 27.4 32.1 32.2 31.5
en 55.6 56.7 56.0 56.3 53.2 53.4 53.4 53.2
es 67.9 71.3 71.7 70.4 62.7 63.7 63.5 63.9
fa 54.4 61.3 58.6 59.6 48.1 51.6 51.6 51.9
fi 34.6 43.2 43.0 39.8 21.9 28.4 28.8 25.9
fil 19.0 21.8 20.5 19.4 10.4 12.5 12.3 11.3
fr 74.8 77.6 76.2 76.9 67.0 69.5 69.9 68.0
hi 20.6 26.5 25.9 23.5 10.5 14.6 14.4 12.3
hr 46.9 57.4 56.2 53.1 33.6 41.2 40.4 37.9
hu 47.6 55.2 53.9 53.9 37.4 42.7 43.1 41.2
id 74.1 77.8 77.7 78.4 65.3 68.0 68.2 67.0
it 73.9 79.1 77.5 77.3 67.3 70.4 70.2 69.9
iw 48.2 56.8 54.5 52.3 37.5 43.9 43.1 40.8
ja 48.8 57.8 57.9 53.9 38.1 43.2 42.0 39.6
ko 58.6 65.6 63.3 63.3 50.0 53.4 52.5 51.7
mi 0.7 0.7 0.8 0.6 0.2 0.4 0.3 0.3
nl 61.8 67.8 65.0 66.2 55.1 58.8 59.0 57.7
no 61.0 68.9 67.1 66.8 45.5 52.5 52.0 50.9
pl 62.3 68.9 67.4 67.1 54.0 57.6 59.0 57.1
pt 69.6 71.1 70.1 71.4 61.5 62.6 63.3 63.2
quz 7.2 7.7 7.9 7.7 3.1 3.0 3.2 2.9
ro 52.7 61.4 60.5 60.0 39.8 48.5 48.3 46.0
ru 66.7 70.8 69.5 70.3 61.3 64.2 64.3 64.2
sv 66.8 71.9 71.9 70.1 51.3 55.3 56.6 54.4
sw 9.2 10.2 10.0 9.4 4.6 5.2 5.4 4.4
te 1.1 1.4 1.6 1.0 0.5 0.5 0.7 0.5
th 29.5 38.2 35.1 34.5 21.5 24.7 23.1 24.0
tr 52.9 58.0 57.9 57.6 44.3 47.3 46.9 46.6
uk 52.2 58.1 56.1 56.6 38.7 44.3 44.9 42.8
vi 73.8 79.3 78.4 77.8 61.8 65.7 66.4 65.5
zh 55.3 60.2 60.5 58.7 44.9 48.3 49.8 47.1
Table 5: Per-language retrieval evaluations using the Crossmodal-3600 dataset (Thapliyal et al., 2022).
Appendix C Zero-shot Evaluation on Common Sense Reasoning Tasks
This section describes the evaluation setup for assessing the zero-shot common sense reasoning capabilities of the language model (Table 1).
Hi Ibrahim, I’ve read your code and here’s a detailed description of the evaluation setup for a research paper:
Evaluation Setup for Zero-Shot Common Sense Reasoning
This section describes the evaluation setup for assessing the zero-shot common sense reasoning capabilities of the language model.
Datasets.
Each model is evaluated on four benchmarks:
  • WinoGrande: A multiple-choice dataset focusing on commonsense knowledge (Sakaguchi et al., 2019).
  • OpenBookQA: A multiple-choice question answering dataset requiring knowledge from an open book of elementary level science facts (Mihaylov et al., 2018).
  • BoolQ: A yes/no question answering dataset requiring the model to determine the truthfulness of a given statement based on a provided passage (Clark et al., 2019).
  • PIQA: A multiple-choice benchmark focusing on physics-related reasoning, such as how we interact with objects in daily life (Bisk et al., 2020).
  • SIQA: A multiple-choice dataset focusing on social common sense reasoning (Sap et al., 2019).
Prompting.
Each model is evaluated in a zero-shot setting, meaning it does not receive any training examples specific to the task. Instead, since all tasks above have multiple-choice answers, we evaluate the log-perplexity of each option, applying a causal mask, after concatenating the contextual information (if any), followed by the prefix and the answer. Then, we select the choice that has the lowest per-token log-perplexity score as the model’s answer.
We do not apply any prompting, except in PIQA where we formulate each sentence in the form The goal is: {goal} The solution is: {sol}. We do this in PIQA because, otherwise, the sentences can be difficult to understand, even for humans. In one example, for instance, the goal is: ”Deep clean coffee grinder.” and the two possible solutions are: ”Scrape with rice.” and ”Scrape with flour.” Concatenating directly would result in sentences like ”Deep clean coffee grinder. Scrape with rice.” whose are unclear.
Evaluation Metric.
The evaluation metric is accuracy. For each example, the model’s predictions are compared to the ground truth labels, and the accuracy is calculated as the percentage of correct predictions. The model is evaluated in a deterministic mode, meaning no randomness is involved in the inference process.
Appendix D Recursion and Self-Similarity
Figure 9: Caption
How does depth-wise model recursion relate to self-similarity of language across the temporal dimension? We provide an illustration of this in Figure 9.
Informally, one might view the early layers in a neural network to be transforming the raw linguistic input into a richer language that is appended with useful, high-level details. For instance, in the sentence “the evil queen appears …”, early layers potentially through mechanisms like attention, might modify the token “queen” to incorporate features such as part-of-speech (POS) tagging, sentiment, and grammatical information. Subsequent layers can then iteratively incorporate progressively higher-level contextual information.
The concept of self-similarity in temporal sequences, including language, hints that higher-order linguistic features exhibit patterns analogous to those present in the raw input. If this property holds for language, it implies that the same neural network block can be recursively applied to generate increasingly abstract features.
However, while this interpretation offers a conceptual framework, it is essential to acknowledge that it remains a hypothesis. Empirical evidence has demonstrated the self-similarity of log-perplexity scores in language (i.e. surprise of information-theoretic complexity), indicating that patterns observed at the word level are mirrored at sentence and potentially higher levels of linguistic structure. The efficacy of Recursive INference Scaling (RINS) is another evidence in favor of the informal picture above.
D.1 Architecture Sweep
We use the following pseudocode when sweeping across recursive architectures. The goal is to ensure that we only compare architectures that have a similar size.
# baseline
add(signature=’A’, degree=1)
# RAO
add(signature=’AA’, degree1)
add(signature=’AAA’, degree1)
add(signature=’AAAA’, degree1)
# other architectures
for signature in [’ABB’, ’ABA’, ’AAB’, ’ABBC’, ’AABC’, ’ABCC’, ’ABBB’, ’AAAB’, ’AABB’]:
for degree in [1, 2, 3]:
num_unique_blocks = len(set(signature) ** degree unique_layers
num_layers_per_block = num_layer_in_baseline_model // num_unique_blocks
if num_layers_per_block > 0:
add(signature=signature, degree=degree)
D.2 Training Configuration
D.2.1 Language Modeling
"""Config for training a decoder-only language model."""
config.seed = 0
config.total_steps = ... # swept
config.vocab_size = 32,000
config.input = dict()
config.input.max_len = 1,024 # or 1,536 for long-sequence baseline
config.tokenizer = ’c4_en’
config.input.data = { # equal weight
’c4/en’: 1.0,
’huggingface:cerebras__slimpajama_627b’: 1.0,
for dataset_name in config.input.data:
config.input[dataset_name] = {}
config.input[dataset_name].data = dict(
name=dataset_name,
split=’train’,
config.input[dataset_name].shuffle_buffer_size = 250,000
config.input.batch_size = 1,024
# Optimizer section
config.optax_name = ’scale_by_adam’
config.grad_clip_norm = 1.0
config.lr = 5e-4
config.wd = 5e-5
config.schedule = dict(decay_type=’rsqrt’,
warmup_steps=5,000,
cooldown_steps=5,000,
D.2.2 Vision: Image Classification
"""Config for training a ViT model."""
config.seed = 0
config.total_epochs = ... # swept
config.num_classes = 1,000
config.init_head_bias = -6.9
config.input = dict()
config.input.batch_size = 1,024
config.input.shuffle_buffer_size = 250,000
# preprocessing
config.input.pp = ’value_range(-1, 1)|inception_crop(224)|flip_lr’
config.mixup = dict(p=0.2, fold_in=None)
# Optimizer section
config.optax_name = ’scale_by_adam’
config.grad_clip_norm = 1.0
config.optax = dict(mu_dtype=’bfloat16’, b2=0.95)
config.lr = 0.001
config.wd = 0.0001
config.schedule = dict(decay_type=’cosine’, warmup_steps=10,000)
D.2.3 Language-Image Pretraining
"""Config for training a SigLIP model."""
config.seed = 0
config.total_examples = ... # swept
config.input = dict()
config.input.batch_size = 1,024 x 32
config.input.shuffle_buffer_size = 250,000
# preprocessing
config.input.tokenize = mc4 # multilingual c4
config.input.max_len = 64 # for text
config.input.prefetch = 1 # save host memory
# model
config.model.bias_init = -10.0
config.model.temperature_init = 10.0
# Optimizer section
config.optax_name = ’scale_by_adam’
config.grad_clip_norm = 1.0
config.lr = 0.001
config.wd = 0.0001
config.schedule = dict(decay_type=’cosine’, warmup_steps=20_000)